A certain brand of freezer is advertised to use 880 kW⋅h of energy per year.

P=400 W

1)If the freezer keeps its interior at a temperature of -4.0 ∘C in a 20.0 ∘C room, what is its theoretical maximum performance coefficient?

2)What is the theoretical maximum amount of ice this freezer could make in an hour, starting with water at 20.0

∘C?
Answer #1

1) Kcarnot = Tc/(Th -Tc)

Th = 293 K, Tc = 269 K

Kcarnot = 269/293-269 = 11.21

2) W = Pt = (400 W)(3600 s) =1.44×10^6 J
Qc = K W =1.61×10^7 J.

Qc = mc ΔT + mL

m = 1.61×10^7/(4190*20+334*1000)

m = 38.53 kg.